An instrument shows the number of revolutions per minute made by each tire of a car. In each revolution, the car travels a distance equal to the circumference of one of its tires. The circumference of each tire is equal to 2πr, where r is the radius of the tire.
Maria gets new tires for her car. The radius of each of her old tires is 0.30 meter, and the radius of each of her new tires is 11% larger than the radius of one of her old tires. What is the circumference of each new tire, to the nearest tenth of a meter?

Answer by Alan3354(69443)

An instrument shows the number of revolutions per minute made by each tire of a car. In each revolution, the car travels a distance equal to the circumference of one of its tires. The circumference of each tire is equal to 2πr, where r is the radius of the tire.
Maria gets new tires for her car. The radius of each of her old tires is 0.30 meter, and the radius of each of her new tires is 11% larger than the radius of one of her old tires. What is the circumference of each new tire, to the nearest tenth of a meter?
======================
r of new tires = 0.3*1.11 = 0.333 meters
C = 2*pi*r